OpenStudy (anonymous):

find f(x) and g(x) as y=f(g(x)) please help y=8/(x^2)+2

OpenStudy (anonymous):

\[y=\frac{ 8 }{ x^2 }+2\]

OpenStudy (anonymous):

if you set \[g(x)=x^2\] then \[f(x)={8\over x}+2\]

OpenStudy (anonymous):

you try to find the simplest repeating entity of "x"

OpenStudy (anonymous):

so i would set x^2 at the numerator ?

OpenStudy (anonymous):

for g(x)

OpenStudy (anonymous):

yes g(x) is just x^2

OpenStudy (anonymous):

\[f(g(x))=\frac{8}{g(x)}+2={8\over x^2}+2\]

OpenStudy (anonymous):

oh ok now i see it :) thank you!:D

OpenStudy (anonymous):

yer welcome.:)